Circular No. 7733

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2001 T3 (NEAT) K. Lawrence, S. Pravdo, and E. F. Helin, Jet Propulsion Laboratory,

report the discovery by the NEAT program of a comet with a faint coma on CCD images taken with the Palomar 1.2-m Schmidt telescope. The object also appears cometary on CCD images taken by P. Pravec and P. Kušnirák at Ondřejov (moderately condensed coma of diameter 0′.2 on Oct. 14.9 UT) and by J. Tichá, M. Tichý, and P. Jelínek at Kleť (diffuse 11′′ coma on Oct. 14.9; 10′′ coma and m_{1} = 17.0 on Oct. 15.8).

2001 UT  α_{2000}  δ_{2000}  m_{1}
Oct. 14.44977  1^{h}45^{m}17.^{s}08  −8^{o}27′56′′.7  18.4

The available astrometry (including prediscovery data by LINEAR on Oct. 11) and the following orbital elements appear on MPEC 2001-T58.

T = 2002 Jan. 27.323 TT  ω = 354.^{o}305
e = 0.60172  Ω = 56.739 }2000.0
q = 2.52309 AU  i = 19.107

a = 6.33492 AU n^{o} = 0.061815 P = 15.94 years

2001TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
Oct. 8 1^{h}50.^{m}03 − 8^{o}39.^{′}7 1.727 2.688 159.^{o}9 7.^{o}3 17.5
18 1 42.54 − 8 18.3 1.696 2.661 161.9  6.7 17.4
28 1 34.72 − 7 38.2 1.690 2.636 157.6  8.3 17.3
Nov. 7 1 27.56 − 6 38.0 1.710 2.613 149.6 11.1 17.3
17 1 21.93 − 5 18.4 1.753 2.593 140.4 14.1 17.4
2001 QT_{297}

J. L. Elliot, Massachusetts Institute of Technology, reports on followup observations of ‘Deep Ecliptic Survey Team’ discovery targets as part of an ongoing program initiated by Elliot and S. D. Kern at the 6.5-m Baade telescope at Las Campanas: “Observations of 2001 QT_{297} (MPEC 2001-T38) obtained by D. J. Osip and S. M. Burles using the MagIC instrument on Oct. 11 and 12 UT (with seeing < 0′′.45 on both nights) reveal this to be a binary system. The fainter member of the pair lies ≈ 0′′.6 from the brighter member at p.a. 240^{o}. The brighter member of the pair is only 0.55 magnitude brighter than the companion. Additional photometric and astrometric observations of 2001 QT_{297} are strongly encouraged.”

2001 October 15  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green